# Implementing Cumulative Functions with Generalized Cumulative Constraints

### v1

@@ 9 $\overline{e}_{c}$ 0 0 0
Table 1: Time points computed from the tasks of Example 3.
Figure 3: Representation of the tasks of Example 3 with the resulting minimum ($\underline{P}$) and maximum ($\overline{P}$) profile.
Filtering
@@
$\forall tp\in P\,|\,\underline{s}_{i}\leq tp.time<\min(\overline{s}_{i},\underline{e}_{i}),$
$tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ (4)
$\Rightarrow\underline{s}_{i}\geq tp.next.time$
$\forall tp\in P\,|\,\max(\overline{s}_{i},\underline{e}_{i})\leq tp.time<\overline{e}_{i},$
$tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ (5)
$\Rightarrow\overline{e}_{i}\leq tp.prev.time$
@@
The second rule detects if the execution of a task $i\in T$ at a time point $tp$ is necessary to avoid a violation of the resource capacity range at this time point. If this is the case, the task is adjusted to make sure it is present during the time point: 1) if the task is optional it is set to present; 2) its time window is adjusted such that it executes over the whole time point; 3) its height is adjusted to avoid the profile range violation. In formal terms, the rule is defined as:
$\forall tp\in P\,|\,tp.time\in[\underline{s}_{i},\overline{e}_{i})\wedge tp.\#fp>0,$
$tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C}$
$\Rightarrow\begin{cases}&p_{i}\leftarrow true\\ &\overline{s}_{i}\leq tp.time\\ &\underline{e}_{i}\geq tp.next.time\\ &\underline{c}_{i}\geq deficit\text{ if }deficit>0\\ &\overline{c}_{i}\leq overload\text{ if }overload<0\end{cases}$ (6)
where $deficit=\underline{C}-(\overline{P}-\max(\overline{c}_{i},0))$ and $overload=\overline{C}-(\underline{P}-\min(\underline{c}_{i},0))$ are the deficit and overload of height needed for the profile range to intersect the capacity range.
@@
1). If the task has a fixed part, its height is checked and adjusted directly at each time point of the fixed part:
$\forall tp\in P\mid\overline{s}_{i}\leq tp.time<\underline{e}_{i},$
$\overline{c}_{i}\leq\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge$ (7)
$\underline{c}_{i}\geq\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
@@
2). If the task has no fixed part, the maximum available height over the minimum overlapping interval of the task is used to adjust its height:
$\overline{c}_{i}\leq\max_{\mathclap{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.time\leq\overline{s}_{i}}}\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge$
$\underline{c}_{i}\geq\min_{\mathclap{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.time\leq\overline{s}_{i}}}\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$ (8)
The minimum overlapping interval was introduced by (Gay, Hartert, and Schaus 2015b) and is defined for a task $i\in T$ as the interval $[\underline{e}_{i}-1,\overline{s}_{i}]$. Intuitively, it corresponds to the smallest interval such that, no matter its start time, length or end time, the task executes during at least one time unit of this interval.
@@
The last rule is inspired by (Ouellet and Quimper 2019). It adjusts the maximum duration $\overline{d}$ of the tasks that do not have a fixed part (for the tasks with a fixed part, the adjustment would be redundant with the Forbid rule). The principle is to find the longest time span during which the task can be scheduled without a resource capacity violation.
The function timetabling of Algorithm 3 checks these rules for each task $i\in T$. This algorithm receives as input the set of tasks $T$ and the bounds of the resource capacity $[\underline{C},\overline{C}]$. The algorithm iterates over all non-fixed tasks at Line 3.
1 $P\leftarrow\text{initializeTimeLine}(T)$
2 $O\leftarrow\{i\in T\mid p_{i}=\{true,false\}\}$
3 $R\leftarrow\{i\in T\mid p_{i}=true\}$
@@
6 $tpf\leftarrow tp(\underline{s}_{i})$
7 while $tpf.time<\min(\overline{s}_{i},\underline{e}_{i})$ do
8    if $tpf.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpf.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
9       $\underline{s}_{i}\leftarrow\max(tpf.next.time,\underline{s}_{i})$
@@ 6 $tpf\leftarrow tp(\underline{s}_{i})$
11    else
[…]
13
14    $tpf\leftarrow tpf.next$
@@
16 $tpb\leftarrow tp(\overline{e}_{i}).prev$
17 while $tpb.next.time>\max(\overline{s}_{i},\underline{e}_{i})$ do
18    if $tpb.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpb.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
19       $\overline{e}_{i}\leftarrow\min(tpb.time,\overline{e}_{i})$
20
21    else
[…]
23
24    $tpb\leftarrow tpb.prev$
@@
26 if $\overline{s}_{i}<\underline{e}_{i}$ then
27    while $tpf.time<\underline{e}_{i}$ do
[…]
29       $\underline{c}^{\prime}\leftarrow\underline{C}-(tpf.\overline{P}-\max(\overline{c}_{i},0))$
30       $\overline{c}^{\prime}\leftarrow\overline{C}-(tpf.\underline{P}-\min(\underline{c}_{i},0))$
31       if $i\in R$ then
32          $\underline{c}^{\prime}\leftarrow\underline{c}^{\prime}+\min(\overline{c}_{i},0)$
33          $\overline{c}^{\prime}\leftarrow\overline{c}^{\prime}+\max(\underline{c}_{i},0)$
34
35       $\underline{c}_{i}\leftarrow\max(\underline{c}^{\prime},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{\prime},\overline{c}_{i})$
36       $tpf\leftarrow tpf.next$
37
38 else
39    $\overline{d}^{*}\leftarrow 0$; $s^{\prime}\leftarrow\underline{s}_{i}$
40    $\underline{c}^{*}\leftarrow\underline{C}-tpf.prev.\overline{P}+\max(\overline{c}_{i},0)$
41    $\overline{c}^{*}\leftarrow\overline{C}-tpf.prev.\underline{P}+\min(\underline{c}_{i},0)$
42    while $tpf.time<\overline{s}_{i}$ do
43       $\overline{d}^{*}\leftarrow\max(tpf.time-s^{\prime},\overline{d}^{*})$
44       if $tpf.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpf.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
45          $s^{\prime}\leftarrow tpf.next.time$
46
47       else
[…]
49
50       $\underline{c}^{*}\leftarrow\min(\underline{C}-tpf.\overline{P}+\max(\overline{c}_{i},0),\underline{c}^{*})$
51       $\overline{c}^{*}\leftarrow\max(\overline{C}-tpf.\underline{P}+\min(\underline{c}_{i},0),\overline{c}^{*})$
52       $tpf\leftarrow tpf.next$
53
54    $\overline{d}^{*}\leftarrow\max(\overline{e}_{i}-s^{\prime},\overline{d}^{*})$; $\overline{d}_{i}\leftarrow\min(\overline{d}^{*},\overline{d}_{i})$
55    $\underline{c}_{i}\leftarrow\max(\underline{c}^{*},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{*},\overline{c}_{i})$
56
[…]
Each task is processed in three steps:
First, the Forbid rule is used to adjust the minimum start time of the task by iterating forward over the time points in the interval $[\underline{s}_{i},\min(\overline{s}_{i},\underline{e}_{i}))$ in the loop at Line 3. Second, the Forbid rule is used to adjust the maximum end time of the task by iterating backward over the interval $[\max(\overline{s}_{i},\underline{e}_{i}),\overline{e}_{i})$ in the loop at Line 3.
The last step iterates over the time points of the remaining part of the time window. It depends if the task has a fixed part or not. In the former case, the loop at Line 3 iterates over the time points in the fixed part of the task and adjusts its height following the 1st case of the Height rule. In the case where the task has no fixed part, the loop at Line 3 is executed. The height of the task is adjusted according to the 2nd case of the Height rule and its maximum length is adjusted following the Length rule.
@@ 39    $\overline{d}^{*}\leftarrow 0$; $s^{\prime}\leftarrow\underline{s}_{i}$
Similarly, the minimum and maximum available heights are computed as $\underline{c}^{*}$ and $\overline{c}^{*}$ and used to update the height of the task if it has no fixed part (Line 3). Note that these variables are initialized based on the height available at the time point before the the start of Loop 3 in order to consider the whole minimum overlapping interval of the task.
@@
1 if  #$\#fp>0\wedge(tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C})$  then
2 $p_{i}\leftarrow true$
3 $\overline{s}_{i}\leftarrow\min(tp.time,\overline{s}_{i})$
4 $\underline{e}_{i}\leftarrow\max(tp.next.time,\underline{e}_{i})$
5 $deficit\leftarrow\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
6 $overload\leftarrow\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))$
7 if $deficit>0$ then $\underline{c}_{i}\leftarrow\max(deficit,\underline{c}_{i})$
8 if $overload<0$ then $\overline{c}_{i}\leftarrow\min(overload,\overline{c}_{i})$

### v2

@@ 9 $\overline{e}_{c}$ 0 0 0
Table 1. Time points computed from the tasks of Example 1.
Figure 3 shows a representation of the tasks and the resulting profile ranges. Note that the events $\overline{s}_{c}$ and $\underline{e}_{c}$ are not considered as task $C$ has no fixed part.
Figure 3. Representation of the tasks of Example 1 with the resulting minimum ($\underline{P}$) and maximum ($\overline{P}$) profile.
3.4. Filtering
@@
(4) $\forall tp\in P\,|\,\underline{s}_{i}\leq tp.\mathit{time}<\min(\overline{s}_{i},\underline{e}_{i}),\quad tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}\Rightarrow\underline{s}_{i}\geq tp.\mathit{next}.\mathit{time}$
(5) $\forall\mathit{tp}\in P\,|\,\max(\overline{s}_{i},\underline{e}_{i})\leq\mathit{tp.\mathit{time}}<\overline{e}_{i},\quad tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}\Rightarrow\overline{e}_{i}\leq\mathit{tp.\mathit{prev}.\mathit{time}}$
@@
The second rule detects if the execution of a task $i\in T$ at a time point $tp$ is necessary to avoid a violation of the resource capacity range at this time point. If this is the case, the task is adjusted to make sure it is present during the time point:
  1. (1)
    if the task is optional, it is set to present;
  2. (2)
    its time window is adjusted such that it executes over the whole time point;
  3. (3)
    its height is adjusted to avoid the profile range violation.
In formal terms, the rule is defined as:
$\forall\mathit{tp}\in P\,|\,\mathit{tp.\mathit{time}}\in[\underline{s}_{i},\overline{e}_{i})\wedge\mathit{tp.\mathit{\#fp}}>0,$
(6) $\mathit{tp}.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee\mathit{tp}.\overline{P}-\max(\overline{c}_{i},0)<\underline{C}\Rightarrow\begin{cases}&p_{i}\leftarrow\mathit{true}\\ &\overline{s}_{i}\leq\mathit{tp.\mathit{time}}\\ &\underline{e}_{i}\geq\mathit{tp.\mathit{next}.\mathit{time}}\\ &\underline{c}_{i}\geq\mathit{deficit}\text{ if }\mathit{deficit}>0\\ &\overline{c}_{i}\leq\mathit{overload}\text{ if }\mathit{overload}<0\end{cases}$
where $\mathit{deficit}=\underline{C}-(\overline{P}-\max(\overline{c}_{i},0))$ and $\mathit{overload}=\overline{C}-(\underline{P}-\min(\underline{c}_{i},0))$ are the deficit and overload of height needed for the profile range to intersect the capacity range.
@@
  1. (1)
    If the task has a fixed part, its height is checked and adjusted directly at each time point of the fixed part:
    (7) $\forall tp\in P\mid\overline{s}_{i}\leq tp.\mathit{time}<\underline{e}_{i},\quad\overline{c}_{i}\leq\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge\underline{c}_{i}\geq\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
@@
  2. (2)
    If the task has no fixed part, the maximum available height over the minimum overlapping interval of the task is used to adjust its height:
    (8) $\overline{c}_{i}\leq\max_{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.\mathit{time}\leq\overline{s}_{i}}\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))$
    (9) $\underline{c}_{i}\geq\min_{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.\mathit{time}\leq\overline{s}_{i}}\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
    The minimum overlapping interval was introduced by (gay2015time) and is defined for a task $i\in T$ as the interval $\underline{e}_{i}-1,\overline{s}_{i}]$. Intuitively, it corresponds to the smallest interval such that, no matter its start time, length or end time, the task executes during at least one time unit of this interval.
@@
The last rule is inspired by (ouelletProcessing). It adjusts the maximum duration $\overline{d}$ of the tasks that do not have a fixed part (for the tasks with a fixed part, the adjustment would be redundant with the Forbid rule). The principle is to find the longest time span during which the task can be scheduled without a resource capacity violation. Again, we rely on the internal propagation rules of the interval variables to set the task as absent if the adjustment would empty its domain and to update the $\underline{s}_{i}$ and $\overline{e}_{i}$ attributes following the relations $\underline{s}_{i}\geq\underline{e}_{i}-\overline{d}_{i}$ and $\overline{e}_{i}\leq\overline{s}_{i}+\overline{d}_{i}$. Formally, the rule is written as:
$\overline{d}_{i}\leq\max_{[a,b)\in A}b-a$
(10) where
$A=\left\{[a,b)\subseteq[\underline{s}_{i},\overline{e}_{i}]\mid\forall tp\in P\text{ with }tp.\mathit{time}\in[a,b),\;tp.\underline{P}+\max(\underline{c}_{i},0)\leq\overline{C}\wedge tp.\overline{P}+\min(\overline{c}_{i},0)\geq\underline{C}\right\}$
Intuitively, $A$ is the set of intervals where the task can be scheduled without underloading the minimum and overloading maximum capacity.
Algorithm
The function timetabling of Algorithm 3 enforces these rules for each task $i\in T$. This algorithm receives as input the set of tasks $T$ and the bounds of the resource capacity $\underline{C},\overline{C}]$. The algorithm iterates over all non-fixed tasks at Line 3. Each task is processed in three steps:
First (step 1), the Forbid rule is used to adjust the minimum start time of the task by iterating forward over the time points in the interval $\underline{s}_{i},\min(\overline{s}_{i},\underline{e}_{i}))$ in the loop at Line 3. Second (step 2), the Forbid rule is used to adjust the maximum end time of the task by iterating backward over the interval $\max(\overline{s}_{i},\underline{e}_{i}),\overline{e}_{i})$ in the loop at Line 3.
The last step iterates over the time points of the remaining part of the time window. It depends if the task has a fixed part or not. In the former case (step 3), the loop at Line 3 iterates over the time points in the fixed part of the task and adjusts its height following the 1st case of the Height rule. In the case where the task has no fixed part (step 3’), the loop at Line 3 is executed. The height of the task is adjusted according to the 2nd case of the Height rule and its maximum length is adjusted following the Length rule. These three steps are illustrated in Figure 4.
Figure 4. Timetabling processing order for a task with a fixed part (top) and one without (bottom). Those steps (1, 2, 3, and 3’) are also referenced in Algorithm 3 comments.
$P\leftarrow\text{initializeTimeLine}(T)$
// Compute the profile
$O\leftarrow\{i\in T\mid p_{i}=\{\mathit{true},\mathit{false}\}\}$; $R\leftarrow\{i\in T\mid p_{i}=\mathit{true}\}$
// Optional/Required tasks
@@
 $\mathit{tpf}\leftarrow tp(\underline{s}_{i})$
 // Get the time point at minimum start
3 while $\mathit{tpf}.\mathit{time}<\min(\overline{s}_{i},\underline{e}_{i})$ do // Step 1 (forward): adjust minimum start time
    // Check if task can start at this time point, if not push its start
4    if $\mathit{tpf}.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak\mathit{tpf}.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then $\underline{s}_{i}\leftarrow\max(\mathit{tpf}.\mathit{next}.\mathit{time},\underline{s}_{i})$
5    else $\text{checkIfMandatory}(i,\mathit{tpf},\underline{C},\overline{C})$
    $\mathit{tpf}\leftarrow\mathit{tpf}.\mathit{next}$
    // Move to next time point
6
 $\mathit{tpb}\leftarrow tp(\overline{e}_{i}).\mathit{prev}$
 // Get the time point just before maximum end
7 while $\mathit{tpb}.\mathit{next}.\mathit{time}>\max(\overline{s}_{i},\underline{e}_{i})$ do // Step 2 (backward): adjust maximum end time
    // Check if task can end after this time point, if not pull its end
8    if $\mathit{tpb}.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak\mathit{tpb}.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then $\overline{e}_{i}\leftarrow\min(\mathit{tpb}.\mathit{time},\overline{e}_{i})$
9    else $\text{checkIfMandatory}(i,\mathit{tpb},\underline{C},\overline{C})$
    $\mathit{tpb}\leftarrow\mathit{tpb}.\mathit{prev}$
    // Move to previous time point
@@
11 if $\overline{s}_{i}<\underline{e}_{i}$ then // Step 3: Adjust height (task with a fixed part)
12    while $\mathit{tpf}.\mathit{time}<\underline{e}_{i}$ do // Scan over fixed part $\overline{s}_{i},\underline{e}_{i}-1,]$
13       $\text{checkIfMandatory}(i,\mathit{tpf},\underline{C},\overline{C})$
       $\underline{c}^{\prime}\leftarrow\underline{C}-(\mathit{tpf}.\overline{P}-\max(\overline{c}_{i},0))$; $\overline{c}^{\prime}\leftarrow\overline{C}-(\mathit{tpf}.\underline{P}-\min(\underline{c}_{i},0))$ // Available height range
14       if $i\in R$ then
          $\underline{c}^{\prime}\leftarrow\underline{c}^{\prime}+\min(\overline{c}_{i},0)$; $\overline{c}^{\prime}\leftarrow\overline{c}^{\prime}+\max(\underline{c}_{i},0)$ // Adjust for task’s own contribution P
@@ 13       $\text{checkIfMandatory}(i,\mathit{tpf},\underline{C},\overline{C})$
       $\underline{c}_{i}\leftarrow\max(\underline{c}^{\prime},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{\prime},\overline{c}_{i})$
       // Restrict min/max task consumption
       $\mathit{tpf}\leftarrow\mathit{tpf}.\mathit{next}$
       // Move to next time point
16
17 else // Step 3’: Adjust height and length (task without a fixed part)
    $\overline{d}^{*}\leftarrow 0$
    // Track maximum feasible duration
    $s^{\prime}\leftarrow\underline{s}_{i}$
    // Track start of current feasible interval
    $\underline{c}^{*}\leftarrow\underline{C}-\mathit{tpf}.\mathit{prev}.\overline{P}+\max(\overline{c}_{i},0)$
    // Min height over minimum overlapping interval
    $\overline{c}^{*}\leftarrow\overline{C}-\mathit{tpf}.\mathit{prev}.\underline{P}+\min(\underline{c}_{i},0)$
    // Max height over minimum overlapping interval
18    while $\mathit{tpf}.\mathit{time}<\overline{s}_{i}$ do // Scan over minimum overlapping interval $\underline{e}_{i}-1,\overline{s}_{i}]$
19       $\overline{d}^{*}\leftarrow\max(\mathit{tpf}.\mathit{time}-s^{\prime},\overline{d}^{*})$
20       if $\mathit{tpf}.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak\mathit{tpf}.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then $s^{\prime}\leftarrow\mathit{tpf}.\mathit{next}.\mathit{time}$
21       // Reset startelse $\text{checkIfMandatory}(i,\mathit{tpf},\underline{C},\overline{C})$
22       $\underline{c}^{*}\leftarrow\min(\underline{C}-\mathit{tpf}.\overline{P}+\max(\overline{c}_{i},0),\underline{c}^{*})$
23       $\overline{c}^{*}\leftarrow\max(\overline{C}-\mathit{tpf}.\underline{P}+\min(\underline{c}_{i},0),\overline{c}^{*})$
24       $\mathit{tpf}\leftarrow\mathit{tpf}.\mathit{next}$
@@
    $\overline{d}^{*}\leftarrow\max(\overline{e}_{i}-s^{\prime},\overline{d}^{*})$; $\overline{d}_{i}\leftarrow\min(\overline{d}^{*},\overline{d}_{i})$
    // Apply Length rule
    $\underline{c}_{i}\leftarrow\max(\underline{c}^{*},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{*},\overline{c}_{i})$
    // Restrict min/max task consumption
26
Algorithm 3 timetabling($T,\underline{C},\overline{C}$)
@@
Similarly, the minimum and maximum available heights are computed as $\underline{c}^{*}$ and $\overline{c}^{*}$ and used to update the height of the task if it has no fixed part (Line 3). Note that these variables are initialized based on the height available at the time point before the start of Loop 3 in order to consider the whole minimum overlapping interval of the task.
@@
1if $\mathit{\#fp}>0\wedge(tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C})$ then
2 $p_{i}\leftarrow\mathit{true}$
3 $\overline{s}_{i}\leftarrow\min(tp.\mathit{time},\overline{s}_{i})$
4 $\underline{e}_{i}\leftarrow\max(tp.\mathit{next}.\mathit{time},\underline{e}_{i})$
5 $\mathit{deficit}\leftarrow\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
6 $\mathit{overload}\leftarrow\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))$
7 if $\mathit{deficit}>0$ then $\underline{c}_{i}\leftarrow\max(\mathit{deficit},\underline{c}_{i})$
8 if $\mathit{overload}<0$ then $\overline{c}_{i}\leftarrow\min(\mathit{overload},\overline{c}_{i})$